# Five-loop Anomalous Dimensions of Cubic Scalar Theory from Operator Product Expansion

### v2

@@ 2.1 The cubic scalar theory and its renormalization
in which the symbol X can represent the fields $\varphi,\sigma$, the couplings $h,g$ or the operator $\varphi^{Q}$ of arbitrary value $Q$. The $Z_{{\tiny\texttt{X}}}$-factor can be obtained by determining the coefficients $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ up to the appropriate loop orders. All such $Z$-factors can be conveniently computed using two-point propagator-type integrals, following the spirit of the OPE method to renormalization computations. We will review the OPE method in the next subsection.
2.2 The OPE method for the renormalization
In general quantum field theory, the product of two time-ordered operators admits an asymptotic expansion in the limit where their coordinates coincide, known as the Operator Product Expansion. The expansion basis is a set of composite operators of the theory, ordered according to their dimensions, while the expansion coefficients are functions of the coordinate difference, which are in general singular. When considering the OPE in momentum space by Fourier transformation with respect to the coordinate difference, the expansion is performed in the limit of large momentum $p\to\infty$, schematically expressed as,
@@ 2.2 The OPE method for the renormalization
where $k$ is a finite momentum, and the momenta of the two operators tend to the opposite directions of infinite limits [^1] Pictorially, we can consider the two operators with large momentum as two hard fields propagating in the background of soft fields. This idea has also been realized in the study of the boundary operators of BCFW recursion relations Huang:2022qnx . By shifting two external momenta $p_{1},p_{2}$ with a complex auxiliary momentum $zq$ as $p_{1}\to p_{1}+zq$, $p_{2}\to p_{2}-zq$, at the large $z$ limit we get two hard fields propagating in the background of soft fields.. The superscript $^{\tiny\mbox{ren}}$ in the operator $\widehat{O}$ represents a renormalized operator in dimensional regularization scheme.
Alternatively, we may consider the OPE of bare operators, schematically written as,
@@
Special attention should be paid to the OPE of fundamental fields, since the $Z$-factor of a fundamental field $\varphi$ or $\sigma$ does not correspond directly to $Z_{\varphi}$ or $Z_{\sigma}$. Instead, it is defined through $\varphi=Z_{\varphi}^{-\frac{1}{2}}\varphi_{0}$ or $\sigma=Z_{\sigma}^{-\frac{1}{2}}\sigma_{0}$. Furthermore, one usually computes the amputated correlation functions, and an extra factor $Z_{\varphi}$ or $Z_{\sigma}$ should be included for each fundamental field in the OPE of amputated correlation functions. With these two considerations, a fundamental field $\varphi$ or $\sigma$ contributes a $Z^{\frac{1}{2}}_{\varphi}$ or $Z^{\frac{1}{2}}_{\sigma}$ factor to the numerator of the Wilson coefficient relation (17) in the context of amputated OPE. However, the $Z$-factor in the denominator corresponds to an operator but not an external leg, so it still contributes to the $Z^{-\frac{1}{2}}_{\varphi}$ or $Z^{-\frac{1}{2}}_{\sigma}$ factor.
Due to the fact that the Wilson coefficients of renormalized operators $C_{i}^{\tiny\mbox{ren}}(g_{i};p)$ are UV finite, once the Wilson coefficients of bare operators $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed, the ratios of $Z$-factors can be determined using UV finiteness conditions generated from the cancellation of $\epsilon$ poles. It could be possible that some operators in the basis have the same dimension, leading to the operator mixing situation. In this case, the relation (16) is generalized to the vectors of mixed operators related by matrices of $Z$-factors. The $Z$-matrices can be determined by the UV finiteness conditions from all the OPE coefficients corresponding to the mixed operators. Importantly, the above discussion is based on general operators, which means that it is applicable to either simple operators, or operators with Lorentz indices, with fermion and gauge fields, etc.
The Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ of operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ can be computed by the large momentum expansion of bare correlation functions. For an operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ composed of $n$ fields, the corresponding Wilson coefficient is given by
@@
where $k=-\sum_{i=1}^{n}k_{i}$. Pictorially, this corresponding to splitting the Feynman graph of the original correlation function into hard and soft sub-graphs by cutting $n$ internal propagators and/or external legs, with the Wilson coefficient receiving contributions from all possible $n$-cut hard sub-graphs. If the operator contains fields with derivatives, for instance $\partial_{\mu}\varphi_{0}(k)$, we can compute the coefficient by acting $i\frac{\partial}{\partial k^{\mu}}$ to the bare correlation function. For hard sub-graphs, since all the $k_{i},i=1,\ldots,n$ have been set to zero and there remains only two operators with hard momenta, we effectively get two-point graphs. Hence the Wilson coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed by the two-point propagator-type integrals. This greatly simplifies the computation, especially when dealing with correlation functions with large number of external legs.
The above discussion outlines the core idea of computing renormalization functions via OPE method. To determine beta functions and anomalous dimensions, we first need the $Z$-factors, which are extracted from UV finiteness conditions generated from (17), i.e.,
@@
To establish these UV finiteness conditions, we first need to compute the coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ from the corresponding two-point integrals. Then using the general ansatz (13) for $Z$-factors, we can expand the complete expression as series expansion of $\epsilon$, and generate equations by setting coefficients of all $\epsilon$ poles to zero. The $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ coefficients of the $Z_{\tiny\texttt{X}}$-factor can be obtained by solving these equations, and consequently the $Z_{\tiny\texttt{X}}$-factor can be successfully determined to the desired loop order, provided the two-point integrals can be computed to that order.
Now let us apply the OPE method to the renormalization computation of the cubic scalar theory, working in coordinate space. As in the standard field theory computation, the $Z_{\varphi}$ and $Z_{\sigma}$-factors are computed from two-point correlation functions. This computation can be naturally framed within the OPE method. Consider the OPE,
@@
where $\mathbf{1}$ is the unit operator of zero dimension, and $c$ generically represents Wilson coefficient, which is in general different between the two relations. Since they are the OPE of two fundamental fields, the large momentum expansion of them produces,
@@
where we use notation $G(g_{i0};\varphi(x_{1}),\varphi(x_{2}),\varphi(\mbox{\textst{$x$}}_{3}),\ldots)$ to represent the bare Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$. It is computed by the two-point integrals, derived from the bare correlation function with $x$ denoting the corresponding external leg removed, as in eqn.(18). Since they are already two-point functions, no additional legs need removal, so $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}))$ are computed directly from Feynman graphs of two-point correlation functions. Then from UV finiteness conditions of (21), the $Z$-factors of fields $\varphi$ and $\sigma$ can be determined.
The $Z$-factors of coupling constants $g,h$ can be computed from the OPE,
@@
In the large momentum expansion, it produces $Z$-factor ratios $Z^{\frac{1}{2}}_{\varphi}Z^{\frac{1}{2}}_{\varphi}/Z^{-\frac{1}{2}}_{\sigma}$ and $Z^{\frac{1}{2}}_{\sigma}Z^{\frac{1}{2}}_{\sigma}/Z^{-\frac{1}{2}}_{\sigma}$ respectively. So we obtain,
@@
Further using the relation (7) we get [^2] In the previous paper Huang:2024hsn , we have taken the convention that the $g_{0}$ or $h_{0}$ in the denominator is absorbed into the bare correlation function for three-point functions. This leads to slightly different expressions in that paper, where $G$ represents the bare correlation function divided by $g_{0}$ or $h_{0}$ respectively.
@@
The $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ are directly computed from the two-point integrals of corresponding three-point functions by removing one leg of $\sigma$ field, and the $Z_{g},Z_{h}$ can be determined from UV finiteness conditions of (25), (26). With results of $Z_{\varphi},Z_{\sigma}$ and $Z_{g},Z_{h}$, the beta functions and anomalous dimensions of fields can be readily computed by definitions (9), (10).
The $Z$-factor of the $\varphi^{Q}$ operator can be computed from the OPE,
$\varphi^{Q}\bar{\varphi}~\sim~c_{i}~\varphi^{Q-1}~.$ (27)
After large momentum expansion it produces the ratio of $Z$-factors as,
@@
Thus, the $Z_{\varphi^{Q}}$-factor can be computed recursively using the UV finiteness conditions,
@@
where $G_{Q}$ is computed from Feynman graphs of the $\varphi^{Q}$-$Q\varphi$ correlation function by removing (Q1)(Q-1) external legs. They are two-point integrals with large momentum localized at the $\varphi^{Q}$ operator and one remaining external leg. Using the results of $Z_{\varphi^{Q}}$ and the previously computed beta functions, the anomalous dimension of the $\varphi^{Q}$ operator can be computed by definition (11).
@@ 2.3 From Feynman diagrams to two-point integrals
The UV finiteness conditions discussed in the previous subsection require computing two-point integrals derived from two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions. In this subsection, we review some techniques which will help to simplify the problem of generating Feynman diagrams and transfer them to two-point graphs for subsequent evaluation of two-point integrals.
@@
The cubic scalar theory we consider involves two types of fields $\varphi,\sigma$, and two types of cubic interactions. The complex field $\varphi$ carries a scalar charge that flows along the line of $\varphi,\bar{\varphi}$ fields, which can be represented by an arrow indicating the direction of charge flow. When generating Feynman diagrams, we need to account for all possible field configurations in the internal lines, which becomes a rather difficult task when the number of internal lines is large. It has also been observed that for a given external field configuration, different internal field configurations only contribute to varying combinations of couplings, while the underlying integrals remain identical. Thus there is no need to generate all Feynman diagrams and redundantly compute the same integrals. To clarify the separation between the structure of couplings and the integrals, we will adopt the $\Phi$-field representation. In this framework, we combine the $\varphi_{i},i=1,\ldots,N$ fields and the $\sigma$ field into a single $(N+1)$ component scalar field as,
@@
and the cubic interactions can be universally encoded as
@@
with all other components of the symmetric coupling $\lambda_{IJK}$ vanishing. In the $\Phi$-field representation, we essentially generate the topologies of Feynman diagrams without specifying the explicit configuration of component fields. In this case, we have separated the information of the scalar integral part from the factor of couplings. The Feynman scalar integral can be derived from a given topology by assigning propagators according to Feynman rules, while the factors of couplings of all Feynman diagrams sharing the same topology are encoded in the combination of $\lambda_{IJK}$ couplings, which we define as the coupling factor.
@@
When writing the Feynman integral of correlation functions, we first generate all corresponding topologies with private Mathematica code. For each topology, we get the scalar integral from the combination of propagators, multiplied by the standard symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ from the total symmetry of the graph and external legs. Each topology effectively represents a collection of numerous Feynman diagrams that share the same external state configuration but differ in their internal state configurations. By summing over the indices of the $\lambda_{IJK}$, we get the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of the contributing Feynman graphs.
To clarify the concept of the coupling factor, let us take the two-point two-loop topologies $\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ depicted in Fig.(2) as an example.
@@
For this topology, we can consider the external field configuration of two external $\sigma$ field represented by the dashed lines, or the field configuration of two external $\varphi,\bar{\varphi}$ fields represented by the arrowed lines, as depicted in the LHS of Fig.(1). Each vertex is assigned a $\lambda_{IJK}$, and after summing over the internal indices, we get the respective coupling factors as,
$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\varphi\varphi})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{iI_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}i}=2g^{3}h+2g^{4}~.$ (33)
Alternatively, we can draw all contributing Feynman diagrams with the explicit internal field configurations, as shown in the RHS of Fig.(1). It can be verified that summing over all four Feynman diagrams of each external field configuration indeed produces the coupling factors evaluated above. Since generating topologies is significantly more efficient than generating complete Feynman diagrams, and the evaluation of coupling factors involves only algebraic operations, the $\Phi$-field representation substantially enhances the computational capability.
@@
In the UV finiteness conditions, the function $G$ is computed schematically as
@@
where the summation runs over all possible topologies, and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{{\tiny\mbox{topo}}})$ is computed from the two-point integrals generated by these topologies. Remark that the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of a topology depends on the explicit external field configuration, but the symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}$ are irrelevant to the external fields, they are computed from the topologies. For two-point correlation functions, $G$ is directly evaluated using all 1PI two-point topologies. For three-point correlation functions, we start from all 1PI three-point topologies and compute the independent two-point graphs by removing one external leg from these topologies in all possible ways. For the $\varphi^{Q}$-$Q\varphi$ correlation function, we start from all 1PI $\varphi^{Q}\to Q\varphi$ topologies, and $G_{Q}$ can be computed from all independent two-point graphs generated by removing (Q1)(Q-1) external legs from these topologies in all possible ways.
By investigating the structure of 1PI $\varphi^{Q}\to Q\varphi$ topologies, it has been proven that only a part of these topologies contributes to the two-point integrals Huang:2024hsn . These contributing topologies are called irreducible graphs, characterized by the property that when the operator vertex is deleted, all remaining external legs stay connected in a single graph (except those that reduce to a single line). The other topologies are called reducible graphs, with the property that when deleting the operator vertex causes the external legs to distribute into two or more separate connected components. Hence we only need to generate the irreducible graphs of the 1PI $\varphi^{Q}\to Q\varphi$ topologies, and compute $G_{Q}$ from their corresponding two-point graphs after all possible removals of (Q1)(Q-1) legs. The summation of these graphs gives,
@@
The first term comes from the tree-level contribution and evaluates to 1, while
@@
The $G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}$ receives contributions from all $n_{\ell}$-loop irreducible topologies. As in the standard procedure, it can be evaluated using the symmetry factor, coupling factor and UV divergence as,
@@
where the summation is over all non-isomorphic graphs generated by removing (Q1)(Q-1) external legs of the irreducible 1PI topologies of the $\varphi^{Q}$-$Q\varphi$ correlation function.
Let us summarize this section. After generating the independent topologies of $\varphi^{Q}$-$Q\varphi$ correlation functions, one computes the symmetry factor and coupling factor of each topology, and the UV divergences of two-point integrals in the dimensional regularization scheme to an appropriate $\epsilon$ order. Then the $G$ and $G_{Q}$ can be readily computed by summing over all two-point graphs. The $Z$-factors could be determined from UV finiteness conditions, and eventually the anomalous dimensions can be computed from these $Z$-factors.

### v4

@@ 2.1 The cubic scalar theory and its renormalization
where X represents the fields $\varphi,\sigma$, the couplings $h,g$ or the operator $\varphi^{Q}$. The $Z_{{\tiny\texttt{X}}}$-factor is obtained by determining the coefficients $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ up to the relevant loop orders. All such $Z$-factors can be conveniently computed using two-point propagator-type integrals, following the spirit of the OPE based renormalization algorithm.
2.2 The OPE based algorithm for renormalization
The OPE is a classical technique for investigating physical quantities in the asymptotic expansion limit, and it is textbook knowledge that the anomalous dimension of an operator is linked to the Wilson coefficients of the OPE Collins_1984 . In principle, this connection enables the computation of anomalous dimensions directly from Wilson coefficients. In Huang:2024hsn , we proposed an algorithm that explicitly realize this connection by relating renormalization $Z$-factors to the UV finiteness of Wilson coefficients. This algorithm was validated in practical renormalization calculations for the cubic scalar theory and Gross-Neveu-Yukawa theory, confirming its versatility and efficiency. Unlike other renormalization computational techniques, such as the $R^{\ast}$-operation method, this algorithm treats UV divergences globally in renormalization procedures, eliminating the need to subtract sub-divergences arising from certain Feynman diagrams. It also provides a systematic approach to extracting the essential UV divergences from two-point integrals, applicable not only to beta functions and the anomalous dimensions of fundamental fields, but also to the anomalous dimensions of composite operators, including those involving correlation functions with a large number of external legs.
To describe this algorithm, we begin with a general OPE. The product of two time-ordered operators is asymptotically expanded into a basis of composite operators, where the expansion basis is ordered by the operators’ dimensions. In momentum space, the expansion is performed in the large momentum limit $p\to\infty$, schematically expressed as
@@ 2.2 The OPE based algorithm for renormalization
where $k$ is a finite momentum, and the momenta of the two operators approach infinity in opposite directions [^1] Pictorially, we can regard the two large momentum operators as hard fields propagating in a background of soft fields. This idea has also been implemented in the study of boundary operators in BCFW recursion relations Huang:2016bmv ; Huang:2022qnx . By shifting two external momenta $p_{1},p_{2}$ with a complex auxiliary momentum $zq$ as $p_{1}\to p_{1}+zq$, $p_{2}\to p_{2}-zq$, at the large $z$ limit we get two hard fields propagating in the background of soft fields.. The superscript $^{\tiny\mbox{ren}}$ denotes a renormalized operator within the dimensional regularization scheme. Alternatively, we may consider the OPE of bare operators, schematically written as
@@
This relation explicitly realize the connection between anomalous dimensions and Wilson coefficients. By carefully selecting the basis operator $\widehat{O}_{i}$, we can compute the desired $Z$-factors using simpler correlation functions.
Special attention must be paid to the OPE of fundamental fields. For a fundamental field $\phi$, its associated $Z$-factor does not directly correspond to $Z_{\phi}$. Instead, the field renormalization is defined as $\phi=Z_{\phi}^{-\frac{1}{2}}\phi_{0}$. Furthermore, amputated correlation functions are typically computed, and an additional factor of $Z_{\phi}$ must be included for each fundamental field in the OPE of amputated correlation functions. Incorporating these two considerations, a fundamental field $\phi$ contributes a factor of $Z^{\frac{1}{2}}_{\phi}$ to the numerator of the Wilson coefficient relation (17) in the context of amputated OPEs.
Since the Wilson coefficients of renormalized operators, $C_{i}^{\tiny\mbox{ren}}(g_{i};p)$, are UV finite, once the bare Wilson coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed, the ratios of $Z$-factors can be determined by enforcing UV finiteness conditions, specifically, the cancellation of $\epsilon$ poles. It is possible for basis operators to have identical dimensions, leading to operator mixing. In such cases, the relation (16) generalizes to vectors of mixed operators related by matrices of $Z$-factors. These $Z$-matrices are determined by imposing UV finiteness on all OPE coefficients corresponding to the mixed operators. Notably, the above discussion applies to generic scalar and tensor operators within the framework of general quantum field theories.
The bare Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ for the bare operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ is computed via the large momentum expansion of bare correlation functions. For an operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ composed of $n$ fields, the corresponding Wilson coefficient is given by
@@
where $k=-\sum_{i=1}^{n}k_{i}$. Pictorially, this corresponds to splitting the Feynman graph of the original correlation function into hard and soft sub-graphs by cutting $n$ internal propagators and/or external legs. The Wilson coefficient receives contributions from all possible $n$-cut hard sub-graphs. For operators involving derivative fields, for instance $\partial_{\mu}\varphi_{0}(k)$, the coefficient is computed by acting with $i\frac{\partial}{\partial k^{\mu}}$ on the bare correlation function. For hard sub-graphs, since all $k_{i},i=1,\ldots,n$ are set to zero, leaving only two operators with hard momenta, we effectively reduce the problem to two-point graphs. Thus, the bare Wilson coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are evaluated using two-point propagator-type integrals. This significantly simplifies computations, particularly when dealing with correlation functions with a large number of external legs.
The above discussion outlines the core idea of the OPE based algorithm for computing renormalization functions. To determine beta functions and anomalous dimensions, we first extract $Z$-factors from UV finiteness conditions implied by eqn.(17),
@@
To establish these UV finiteness conditions, firstly we compute the bare coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ from the corresponding two-point integrals. Then using the general ansatz (13) for $Z$-factors, we expand the full expression as a series in $\epsilon$, and generate equations by setting the coefficients of all $\epsilon$ poles to zero. Solving these equations yields the coefficients $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ of the $Z_{\tiny\texttt{X}}$-factor, thereby determining $Z_{\tiny\texttt{X}}$-factor to the desired loop order, provided the two-point integrals can be computed to that order.
We now apply this algorithm to the cubic scalar theory of interest, working in coordinate space. As standard, the $Z_{\varphi}$ and $Z_{\sigma}$-factors are computed from two-point correlation functions. This computation naturally fits within this algorithmic framework. Consider the OPEs,
@@
where $\mathbf{1}$ denotes the zero-dimensional unit operator, and $c,c^{\prime}$ are distinct Wilson coefficients for the two expansions. For the OPE of two fundamental fields, the large momentum expansion yields
@@
where we use the notation $G(g_{i0};\varphi(x_{1}),\varphi(x_{2}),\varphi(\mbox{\textst{$x$}}_{3}),\ldots)$ to represent the bare Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$. As defined in eqn.(18), this coefficient is computed via two-point integrals derived from bare correlation functions, with $x$ indicating the removal of the corresponding external leg. Since they are already two-point functions, no additional legs need ro be removal. Thus $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}))$ are evaluated directly from the Feynman graphs of the two-point correlation functions. The $Z$-factors $Z_{\varphi}$ and $Z_{\sigma}$ are then determined by imposing UV finiteness conditions of (21).
The $Z$-factors for the coupling constants $g$ and $h$ are computed from the OPEs,
@@
In the large momentum limit, these expansions yield $Z$-factor ratios $Z^{\frac{1}{2}}_{\varphi}Z^{\frac{1}{2}}_{\varphi}/Z^{-\frac{1}{2}}_{\sigma}$ and $Z^{\frac{1}{2}}_{\sigma}Z^{\frac{1}{2}}_{\sigma}/Z^{-\frac{1}{2}}_{\sigma}$, respectively. Here, the soft $\sigma$ field is treated as an operator and dose not require amputation. This leads to
@@
Using the coupling renormalization relations (7), we further obtain [^2] In the previous paper Huang:2024hsn , we adopted the convention that $g_{0}$ or $h_{0}$ in the denominator is absorbed into the bare correlation function for three-point functions. This leads to slightly different expressions in that paper, where $G$ represents the bare correlation function divided by $g_{0}$ or $h_{0}$, respectively.
@@
The coefficients $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ are computed directly from two-point integrals of the corresponding three-point functions, with one $\sigma$ field leg removed. The factors $Z_{g}$ and $Z_{h}$ are then determined by enforcing UV finiteness on (25) and (26). With $Z_{\varphi},Z_{\sigma}$ and $Z_{g},Z_{h}$ in hand, the beta functions and field anomalous dimensions are readily computed via definitions (9) and (10).
Finally, the $Z$-factor of the $\varphi^{Q}$ operator is computed from the OPE,
$\varphi^{Q}\bar{\varphi}~\sim~c_{Q}~\varphi^{Q-1}~.$ (27)
After large momentum expansion of this OPE yields the $Z$-factors ratio
@@
Thus, $Z_{\varphi^{Q}}$ can be computed recursively by imposing UV finiteness condition,
@@
where $G_{Q}$ is computed from Feynman graphs of the $\varphi^{Q}$-$Q\varphi$ correlation function with (Q1)(Q-1) external legs removed. These integrals are two-point propagator-type integrals, with large momentum localized at the $\varphi^{Q}$ operator and one remaining external leg. Using the resulting $Z_{\varphi^{Q}}$ and the previously computed beta functions, the anomalous dimension of the $\varphi^{Q}$ operator os obtained via eqn.(11).
As a side remark, the transformation of general integrals into two-point integrals is a common technique for evaluating UV divergences across various scenarios. Examples include the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , the IR-rearrangement technique Vladimirov:1979zm ; Chetyrkin:1980pr ; Caswell:1981ek , and works that do not rely on the OPE approach Fei:2014xta ; Jack:2021ziq . In all these cases, two-point integrals serve as a convenient computational tool for extracting UV divergences from Feynman integrals. However, with the aid of the OPE, two-point integrals are encoded as Wilson coefficients, from which anomalous dimensions can be systematically derived.
@@ 2.3 From Feynman diagrams to two-point integrals
The UV finiteness conditions discussed in the preceding subsection require computing two-point integrals derived from two-point, three-point, and $\varphi^{Q}$-$Q\varphi$ correlation functions. In this subsection, we review key techniques that facilitate simplifying the generation of Feynman diagrams and transforming them into two-point graphs for the subsequent evaluation of two-point integrals.
@@
The cubic scalar theory under consideration involves two types of fields $\varphi,\sigma$ and two classes of cubic interactions. The complex field $\varphi$ carries a scalar charge that flows along $\varphi$ and $\bar{\varphi}$ lines. This charge flow can be represented by an arrow indicating its direction. Generating Feynman diagrams requires accounting for all possible internal field configurations, a task that becomes increasing cumbersome as the number of internal lines grows. Notably, for a fixed external field configuration, different internal field configurations only affect the combinations of couplings, while the underlying scalar integrals remain identical. Thus, there is no need to generate all Feynman diagrams explicitly or redundantly compute identical integrals. To decouple the coupling structure from the integral calculations, we adopt the $\Phi$-field representation. In this framework, we unify the $\varphi_{i},i=1,\ldots,N$ fields and the $\sigma$ field into a single $(N+1)$ component scalar field
@@
with the cubic interactions universally encoded as
@@
and all other components of the symmetric coupling tensor $\lambda_{IJK}$ vanish. In the $\Phi$-field representation, we generate Feynman diagram topologies without specifying the explicit component fields of $\Phi_{I}$. This effectively separates the scalar integral information from the coupling-dependent factors. The Feynman scalar integral for a given topology is derived by assigning propagators according to standard Feynman rules, while the coupling factors for all diagrams sharing that topology are encapsulated in combinations of the $\lambda_{IJK}$ couplings. We refer to this combined term as the coupling factor.
@@
When constructing the Feynman integrals for correlation functions, we first generate all corresponding topologies using custom Mathematica code. For each topology, the scalar integral is derived from the product of propagators, multiplied by the standard symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$, accounting for the total symmetry of the graph and external legs. Each topology effectively represents a set of numerous Feynman diagrams that share the same external state configuration but differ in their internal state configurations. By summing over the indices of the $\lambda_{IJK}$, we obtain the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ for the contributing Feynman graphs.
To illustrate the concept of the coupling factor, we consider the two-point two-loop topology shown below,
For this topology, we analyze two external field configurations: (1) two external $\sigma$ field represented by dashed lines, (2) two external $\varphi_{i}$ and $\bar{\varphi}_{j}$ fields represented by arrowed lines, as depicted in the LHS of Fig.(1). Each vertex is assigned a $\lambda_{IJK}$ tensor, and summing over the internal indices yields the respective coupling factors
$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\varphi_{i}\bar{\varphi}_{j}})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{iI_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}j}=(2g^{3}h+2g^{4})\delta_{ij}~.$ (33)
Alternatively, we can draw all contributing Feynman diagrams with their internal field configurations, as shown in the RHS of Fig.(1). It can be verified that summing over the contributions of all four Feynman diagrams for each external field configuration reproduces the coupling factors derived above. Since generating topologies is significantly more efficient than generating full Feynman diagrams, and evaluating coupling factors involves only algebraic operations, the $\Phi$-field representation substantially enhances computational efficiency.
@@
In the UV finiteness conditions, the function $G$ is schematically computed as
@@
where the summation runs over all possible topologies, and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{{\tiny\mbox{topo}}})$ is evaluated from the two-point integrals generated by these topologies. It is worth noting that the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of a topology depends on the explicit external field configuration, whereas the symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ and UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}$ are topology-dependent quantities that are independent of external fields. For two-point correlation functions, $G$ is directly evaluated using all 1PI two-point topologies. For three-point correlation functions, we start from all 1PI three-point topologies and construct independent two-point graphs by removing one external leg from these topologies in every possible manner. For the $\varphi^{Q}$-$Q\varphi$ correlation function, we begin with all 1PI $\varphi^{Q}\to Q\varphi$ topologies. $G_{Q}$ is then computed from all independent two-point graphs generated by removing (Q1)(Q-1) external legs from these topologies in all possible ways.
By analyzing the structure of 1PI $\varphi^{Q}\to Q\varphi$ topologies, it has been proven that only a subset of these topologies contributes to the two-point integrals Huang:2024hsn . These contributing topologies are termed irreducible graphs, defined by the property that deleting the operator vertex leaves all remaining external legs connected within a single graph (with the exception of configurations that reduce to a single line). The remaining topologies are classified as reducible graphs, which exhibit the property that deleting the operator vertex partitions the external legs into two or more disjoint connected components. Thus, we only need to generate irreducible graphs from the set of 1PI $\varphi^{Q}\to Q\varphi$ topologies, and compute $G_{Q}$ from their corresponding two-point graphs after all possible removals of (Q1)(Q-1) legs. The summation of these graphs yields
@@
The first term corresponds to the tree-level contribution and evaluates to 1, while
@@
$G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}$ receives contributions from all $n_{\ell}$-loop irreducible topologies. Following standard procedures, it can be evaluated using the symmetry factor, coupling factor and UV divergence as
@@
where the summation is performed over all non-isomorphic graphs generated by removing (Q1)(Q-1) external legs from the irreducible 1PI topologies of the $\varphi^{Q}$-$Q\varphi$ correlation function.
Let us summarize this section as follows. After generating the independent topologies of $\varphi^{Q}$-$Q\varphi$ correlation functions, we compute the symmetry factor and coupling factor of each topology, along with the UV divergences of two-point integrals in the dimensional regularization scheme up to the desired $\epsilon$ order. The quantities $G$ and $G_{Q}$ can then be readily obtained by summing over all two-point graphs. The $Z$-factors are determined from the UV finiteness conditions, and the anomalous dimensions are ultimately computed from these $Z$-factors.